• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 00:09:00 on May 7, 2000

    Entry number 40646

    keyword=Beam Time Accounting
    Hall A

    5/6/2000 1600-2400
    Experiment #: E93-049
    Shift Leader: D.W.Higinbotham
    Shift Workers: V. Punjabi, J. Kellie

    16001700180019002000210022002300Total Time
    Acceptable Beam
    in Use
    1 1 50 1 1 55 1 50 7.58333333333333
    Beam Available
    but Not in Use
    0 0 10 0 0 5 0 10 0.416666666666667
    Beam Not Available
    or Unacceptable
    0 0 0 0 0 0 0 0 0
    Accelerator Con-
    figuration Chage
    0 0 0 0 0 0 0 0 0
    Experiment
    Ready
    1 1 50 1 1 55 1 1 7.75
    Planned Config-
    uration Change
    0 0 0 0 0 5 0 0 0.0833333333333333
    Unplanned
    Experiment Down
    0 0 10 0 0 0 0 00.166666666666667
    Hall not
    ready reasons
    Target 0 0 10 0 0 0 0 0 0.166666666666667
    HRSe 0 0 0 0 0 0 0 0 0
    HRSh 0 0 0 0 0 0 0 0 0
    Data Acq 0 0 0 0 0 0 0 0 0
    Other 0 0 0 0 0 0 0 0 0

    Comments: 18:22 Carbon Pointing File#2082 18:40 Problem with the Cryotarget software which made it look like we couldn't move the target ladder. But after IOC reboot everything was fine. 18:50 We start taking data again 19:00 10cm Dummy Data Taken File #2083 20:00 Gas Alarm due to Argon. Jack came in and fixed the problem. 21:42 File 2086 is the BPM pedestal run

    Data Entry by: D.W.Higinbotham